OpenStudy (anonymous):

To factor the trinomial by grouping, you need to rewrite 3x as which expression? 3x^2+3x-36

OpenStudy (anonymous):

\[ 3x^2+3x-36 =3(x^2+ x-12)=3( x+4)(x-3) \]